# Solving Modular Linear Systems with a Constraint by parallel decomposition of the Smith form and extended Euclidean division modulo powers of primes divisors

### v3

@@ Abstract
Integral linear systems $Ax=b$ with matrices $A$, $b$ and solutions $x$ are also required to be in integers, can be solved using invariant factors of $A$ (by computing the Smith Canonical Form of $A$). This paper explores a new problem which arises in applications, that of obtaining conditions for solving the Modular Linear System $Ax=b\mbox{ \rm{mod} }n$ given $A,b$ in $\mathbb{Z}_{n}$ for $x$ in $\mathbb{Z}_{n}$ along with the constraint that the value of the linear function $\phi(x)=\langle w,x\rangle$ is coprime to $n$ for some solution $x$. In this paper we develop decomposition of the system to coprime moduli $p^{r(p)}$ which are divisors of $n$ and show how such a decomposition simplifies the computation of Smith form. This extends the well known index calculus method of computing the discrete logarithm where the moduli over which the linear system is reduced were assumed to be prime (to solve the reduced systems over prime fields) to the case when the factors of the modulus are prime powers $p^{r(p)}$. It is shown how this problem can be addressed effciently using the invariant factors and Smith form of the augmented matrix $[A,-p^{r(p)}I]$ and conditions modulo $p$ satisfied by $w$, where $p^{r(p)}$ vary over all divisors of $n$ with $p$ prime.
@@ 1 Introduction
arises in many applications of Number Theory where, $n$ is a given modulus and $A,b$ are matrices over the modular ring $\mathbb{Z}_{n}$, such as the prime factorization by quadratic residues and the index calculus method for discrete log computation [1, 9] over finite fields. In both these problems main practical hurdles are, the large size of the system (or the number of unknowns) along with the large size of the modulus $n$. For solving such large sized linear modular systems over large $n$ it would be highly desirable to reduce the problem for parallel computation using the Chinese remainder theorem (CRT) over moduli $p^{r(p)}$ which are coprime divisors of $n$. Surprisingly, such a reduction does not seem to have been discussed in the past literature for solving modular linear systems. A well known case of reduction arises for the Baby Step Giant Step method of solving the discrete log over a cyclic group with known prime factorization of the order. This computation of the discrete log over a group of large order $g$ has been shown to be decomposable over groups of order $p$, the prime divisors of $g$, well known as Pohlig-Hellman reduction [2]. On the other hand, reduction of the linear system arising in the index calculus method of computing the discrete logarithm has not been known when the group order has prime power factors. The reduction in the number of variables is usually achieved by special number field sieves [11]. In most discrete log computations using index calculus the prime is of the form $12q+$ 1 where $q$ is a large prime. Hence the linear algebra step (which leads to solving a modular linear system) of solving the logarithms of the factor base by a modular linear system is performed for modulo $q$ and $2$ and the result is computed by CRT. But for general modulus $n$ the reduction of the linear system requires solving for moduli $p^{r}$. In this paper we develop a reduction of the Smith form computation of an extended matrix for solving the modular system over moduli $p^{r}$ which gives the parallel computation necessary for solving the system.
@@
In this paper we consider an extended problem of solving (1) with the constraint that a given linear function $\phi(x)=\langle w,x\rangle$ is coprime to $n$. Such a problem arises in Cryptanalysis, [13, 14], in the problem of inversion of sequences of $m$-tuples over Cartesian products of binary or small prime $p$ fields, $\mathbb{F}_{p}^{m}$ or that of local inversion of non-linear maps over $\mathbb{F}_{p}^{m}$ when the complexity of the sequences is considered in terms of Linear Recurrence Relations (LRR) modulo $p^{m}$. Previous cases of LRR as well as non-linear RRs over the finite field $\mathbb{F}$ are discussed in [12, 13, 14]. For instance when the sequence $\{s_{k},k=0,1,2,\ldots\}$ has minimal linear recurrence, the unknown co-efficients of the minimal polynomial $\{x_{0},\ldots,x_{(m-1)}\}$ satisfy a linear system defined by the Hankel matrix of the sequence an the invertibility of the sequence requires that $x_{0}\neq 0$ which is a special case of a linear function $\langle w,x\rangle=x_{0}\neq 0$ over the field of the sequence. When the LRR is considered over modulo $n$ then the corresponding condition becomes $x_{0}$ invertible modulo $n$. Hence the problem considered in this paper is a general problem of solving the system (1) with the constraint that a given linear function $\langle w,x\rangle$ is coprime to $n$.
@@
There are two main hurdles with this problem of computing LRRs of sequences modulo $p^{r}$. First one is the size of the LRR over $\mathbb{F}_{p}$ which grows rapidly with $m$ comparable to the exponential size $p^{m}$. The second theoretical hurdle is that the linear system modulo $p^{r}$ for $r>1$ is no more over a finite field. Hence linear algebra methods (over fields) do not apply. Hence any computational method which can compute the Linear Complexity of such sequences modulo $p^{m}$ has to be scalable for large sizes. The method of decomposition proposed in this paper utilizes special version of Smith form of matrices over $p^{r}$ and is theoretically suitable to achieve the objective of efficient computation.
@@ 1.1 Summery of contribution
Consider a brief summery of contributions made in this paper and the plan of presentation.
@@
    The modular system solution can be decomposed as shown in Theorem 4 in terms of its residues by using the CRT by decomposing the system modulo $p^{r(p)}$ for the set of all prime divisors $p$ of $n$. However the resultant systems modulo $p^{r}$ are also not solvable by linear algebra over fields when there is an $r(p)>1$. Such a system then can be solved by the special Smith form for the augmented matrix $[A,-p^{r}I]$ as shown in Theorem 3.
@@
    Next it is shown in Lemma 3 that the Smith form of $[A,-p^{r}I]$ can be efficiently computed by the solution of the Bezout identity modulo $p^{r}$. This computation involves modulo $p$ divisions only by which the computation of GCD of elements of $A$ and $p^{r}$ is much simpler in computation compared to using Euclidean division to solve the Bezout identity. This is the crux of the contribution of this paper which shows how the direct solution of the Bezout identity modulo $p^{r}$ by $p$-adic arithmetic is more efficient to compute the Smith form for special augmented matrix $[A,-p^{r}I]$ than using the Euclidean division.
@@
    The above technique leads to the methodology of reducing the Smith form computation of $[A,-nI]$ to $[A,-p^{r}I]$ using only modulo $p$ computation. Hence the methodology is a local prime factor decomposition analogous to the decomposition known as Pohlig-Hellman reduction of the Baby Step Giant Step method of computing the discrete logarithm over cyclic groups. Such a reduction has been missing in the literature since the publication of [2].
@@
    This paper shows how use of $p$-adic arithmetic for computation of the transform of the augmented matrix $[A,-p^{r}I]$ to Smith form using the Bezout identity results into an efficient computation compared to the Euclidean division and Smith form computation of $[A,-nI]$ by parallel computation of Smith forms of $[A,-p^{r}I]$. Further, the computation of Smith form of $[A,-p^{r}I]$ is also accomplished efficiently for known primes $p$ using $p$-adic arithmetic. This is a computational insight which reverses the usual computation of Bezout identity for the of $\gcd{a,p^{r}}$ by using Euclidean division. It shows that direct $p$-adic arithmetic requires only modulo $p$ computations while Euclidean division to find this GCD requires division with respect to multiple divisors and dividends.
@@
    Appendix A, treats the case of solving a linear system modulo $q^{r}$ where $q$ is the power $p^{d}$ of a prime denoting the $p$-adic byte representation of numbers modulo $q^{r}$. The resulting algorithm shows the Smith form computation modulo $q^{r}$ can be carried out in efficient byte operations pre-defined in the computer.
@@
Above points briefly summerize contributions of this paper. We focused mainly on the theory begind the computation and have omitted numerical illustrations, which are truly out of scope and would have required a much longer article. This paper is based on the arxiv version [15].
@@ 1.3 Efficient computation by reduction of Smith form modulo $p^{r}$
    Another important issue leading to the difficulty of solving modular liner system modulo $n$ arises when $n$ has divisors of the form $p^{r(p)}$ with $p$ prime and $r(p)>1$. For such cases solving the reduced linear systems $Ax=b\mbox{ \rm{mod} }p$ over prime fields $\mathbb{F}_{p}$ does not lead to a solution of the system modulo $n$. Hence in such cases, efficient (black box) linear algebra method of [3] cannot be used. Such cases of $r(p)>1$ the modular systems have to be necessarily addressed using the Smith form (or invariant factors) of the matrix $[A,-p^{r(p)}I]$. Reduction of Smith form computation over moduli $p^{r}$ using parallel computation (and CRT) has surprisingly not appeared in the past literature.
@@ 2 Solution of the integral linear system in terms of the Smith form
where $f_{i}$ are unique positive integers such that $f_{1}|f_{2}$, $f_{2}|f_{3}$,…,$f_{r-1}|f_{r}$.
@@ 2.1 Solution of the modular linear system
Then we can determine the solutions of the modular system by solving (4) using the Smith form of $[A,-nI]$.
@@
$P[A,-nI]Q=[S,0]$
where $S$ is a $k\times k$ diagonal nonsingular matrix of invariant factors of $[A,-nI]$ and $[S,0]$ is a $k\times(k+l)$ matrix with $l$ zero columns added. Denote $\tilde{b}=Pb$. The system (4) has a solution iff $x_{0}=S^{-1}Pb$ is integral which is equivalent to $f_{i}|\beta_{i}$. For an arbitrary integral $l$ tuple $x_{1}$, denote $\tilde{x}=[x_{0}^{T},x_{1}^{T}]^{T}$. Then
@@
Since the modular system (1) has a solution iff the equivalent system (4) has an integral solution $x,y$, from the theorem (1) it follows that there exist unimodular matrices $P,Q$ which transform $[A,-nI]$ to $[S,0]$ since integral rank $[A,-nI]$ is equal to $k$. The matrix $S$ is the diagonal nonsingular matrix of invariant factors of $[A,-nI]$. The integral system (4) is then transformed to
@@
[…]
@@
[…]
@@
$[PAQ_{0}-nPQ_{2},-nPQ_{3}]\left[\begin{array}[]{l}x_{0}\\ x_{1}\end{array}\right]=Pb.$
@@ 2.2 Solution of the constrained system
Next consider the constraint $\phi(x)=\langle w,x\rangle$ required to be coprime to $n$ for some solution $x$ of (1) or equivalently that of (4). Consider the $(l+k)$ tuple $[w^{T},0^{T}]^{T}$ obtained by adding $k$ tuple of zero rows to $w$. Denote
@@
where $Q$ is the unimodular matrix in the Smith form decomposition of $[A,-nI]$, $w_{0}$ is the first $l$ rows of $\tilde{w}$ and $w_{1}$ is the remaining $k$ rows of $\tilde{w}$. Then we have
@@
Let $\{p_{1},p_{2},\ldots,p_{r}\}$ denote the set of all prime divisors of $n$. Since all solutions of the modular system (1) are given by (6), the solutions which satisfy the constraint $<w,x>$ coprime to $n$ need to be characterized within this set of all solutions. Existence of such solutions satisfying the constraint and choice of $x_{1}$ to construct a solutionare described by the next Theorem.
@@
$\begin{array}[]{lclll}x_{1}&=&x_{1j}&\mbox{ \rm{mod} }p_{j}&\mbox{ such that $w_{1}\mbox{ \rm{mod} }p_{j}\neq 0$}\\ &=&0&\mbox{ \rm{mod} }p_{j}&\mbox{ whenever $\langle w_{0},x_{0}\rangle\mbox{ \rm{mod} }p_{j}\neq 0$}\end{array}$
@@
Theorem above shows how the Problem (1) can be solved by using the Smith form of $[A,-nI]$ and the knowledge of prime divisors of $n$.

### v4

@@ Abstract
Integral linear systems $Ax=b$ with matrices $A$, $b$ and solutions $x$ are also required to be in integers, can be solved using invariant factors of $A$ (by computing the Smith Canonical Form of $A$). This paper explores a new problem which arises in applications, that of obtaining conditions for solving the Modular Linear System $Ax=b\mbox{ \rm{mod} }n$ given $A,b$ in $\mathbb{Z}_{n}$ for $x$ in $\mathbb{Z}_{n}$ along with the constraint that the value of the linear function $\phi(x)=\langle w,x\rangle$ is coprime to $n$ for some solution $x$. In this paper we develop decomposition of the system to coprime moduli $p^{r(p)}$ which are divisors of $n$ and show how such a decomposition simplifies the computation of Smith form. This extends the well known index calculus method of computing the discrete logarithm where the moduli over which the linear system is reduced were assumed to be prime (to solve the reduced systems over prime fields) to the case when the factors of the modulus are prime powers $p^{r(p)}$. It is shown how this problem can be addressed efficiently using the invariant factors and Smith form of the augmented matrix $A,-p^{r(p)}I]$ and conditions modulo $p$ satisfied by $w$, where $p^{r(p)}$ vary over all divisors of $n$ with $p$ prime.
@@ 1 Introduction
arises in many computational problems of Cryptography and Coding where, $n$ is a given modulus and $A,b$ are matrices over the modular ring $\mathbb{Z}_{n}$, such as for instance in the prime factorization of $n$ by quadratic residues and for discrete log computation [1, 9] over finite fields by the index calculus method. In most such problems main practical hurdles are, the large size of the system (or the number of unknowns) along with the large size of the modulus $n$. For solving such large sized linear modular systems over large $n$ it would be highly desirable to reduce the problem by parallel computation using the Chinese remainder theorem (CRT) over moduli $p^{r(p)}$ which are coprime divisors of $n$. Surprisingly, such a reduction does not seem to have been discussed in the past literature for solving modular linear systems. A well known case of reduction to coprime factors of $n$ arises in the Baby Step Giant Step method of solving the discrete log over a cyclic group with known prime factorization of the group order $n$. This computation of the discrete log over a cyclic group of large order $n$ has been shown to be parallely decomposable over groups of order $p$, the prime divisors of $n$, well known as the Pohlig-Hellman reduction [2]. On the other hand, reduction of the linear system arising in the index calculus method of computing the discrete logarithm has not been known when the group order has prime power factors. The reduction in the number of variables is usually achieved by special number field sieves [11] in factorisation problems. In most discrete log computations using index calculus the prime is of the form $2q+1$ where $q$ is a large prime. Hence the linear algebra step (which leads to solving a modular linear system) of solving the logarithms of the factor base by a modular linear system is performed modulo $q$ and modulo $2$ and the result is computed by CRT. Hence the reduced systems are defined over finite prime fields. But for general modulus $n$ the reduction of the linear system requires solving for moduli $p^{r}$ when $r>1$. In this paper we develop a reduction of the Smith form computation of an extended matrix for solving the modular system (1) over moduli $p^{r}$. Hence the algorithm results in parallel decomposition necessary for solving the system in $p$-adic arithmetic.
@@
In fact, in this paper we consider the modular system problem (1) with the further constraint that a given linear function $\phi(x)=\langle w,x\rangle$ is coprime to $n$. Such a problem arises in Cryptanalysis, [13, 14], in the problem of inversion of sequences of $m$-tuples over Cartesian products of binary or small prime $p$ fields, $\mathbb{F}_{p}^{m}$ or that of local inversion of non-linear maps over $\mathbb{F}_{p}^{m}$ when the complexity of the sequences is considered in terms of Linear Recurrence Relations (LRR) modulo $p^{m}$. Previous cases of LRR as well as non-linear RRs over the finite field $\mathbb{F}$ are discussed in [12, 13, 14]. For instance, when the sequence $\{s_{k},k=0,1,2,\ldots\}$ has minimal linear recurrence, the unknown co-efficients of the minimal polynomial $\{x_{0},\ldots,x_{(m-1)}\}$ satisfy a linear system defined by the Hankel matrix of the sequence and the invertibility of the sequence requires that $x_{0}\neq 0$ which is a special case of a linear function $\langle w,x\rangle=x_{0}\neq 0$ over the field of the sequence. When the LRR is considered over modulo $n$ then the corresponding condition becomes $x_{0}$ invertible modulo $n$. Hence the problem considered in this paper is a general problem of solving the system (1) with the constraint that a given linear function $\langle w,x\rangle$ is coprime to $n$.
@@
There are two main hurdles with this problem of computing LRRs of sequences modulo $p^{r}$. First one is that the order of a LRR of a pseudorandom sequence over $\mathbb{F}_{p}$ can be exponential, comparable to $p^{m}$. The second theoretical hurdle is that the linear system modulo $p^{r}$ for $r>1$ required to be solved for computing the LRR is no more over a finite field. Hence linear algebra methods (over fields) do not apply. Hence any computational method which can compute the Linear Complexity of such sequences modulo $p^{r}$ has to be scalable for large sizes of $p^{r}$. The method of decomposition proposed in this paper utilizes special version of Smith form of matrices over $p^{r}$ and is theoretically suitable to achieve the objective of efficient parallel computation using $p$-adic arithmetic.
@@ 1.1 Summery of contribution
Consider a brief summary of contributions made in this paper and the plan of presentation.
@@
    The modular system solution can be decomposed as shown in Theorem 4 in terms of its residues by using the CRT by decomposing the system modulo $p^{r(p)}$ for the set of all prime divisors $p$ of $n$. However the resultant systems modulo $p^{r}$ are also not solvable by linear algebra over fields when there is an $r(p)>1$. Such a system then can be solved by the special Smith form for the augmented matrix $A,-p^{r}I]$ as shown in Theorem 3.
@@
    Next it is shown in Lemma 3 that the Smith form of $A,-p^{r}I]$ can be efficiently computed by the solution of the Bezout identity modulo $p^{r}$. This computation involves modulo $p$ divisions only by which the computation of GCD of elements of $A$ and $p^{r}$ is much simpler in computation compared to using Euclidean division to solve the Bezout identity. This is the crux of the contribution of this paper which shows how the direct solution of the Bezout identity modulo $p^{r}$ by $p$-adic arithmetic is more efficient to compute the Smith form for special augmented matrix $A,-p^{r}I]$ than using the Euclidean division.
@@
    The above technique leads to the methodology of reducing the Smith form computation of [A,nI][A,-nI] to $A,-p^{r}I]$ using only modulo $p$ computation. Hence the methodology is a local prime factor decomposition analogous to the decomposition known as Pohlig-Hellman reduction of the Baby Step Giant Step method of computing the discrete logarithm over cyclic groups. Such a reduction has been missing in the literature since the publication of [2].
@@
    This paper shows how use of $p$-adic arithmetic for computation of the transform of the augmented matrix $A,-p^{r}I]$ to Smith form using the Bezout identity results into an efficient computation compared to the Euclidean division and Smith form computation of [A,nI][A,-nI] by parallel computation of Smith forms of $A,-p^{r}I]$. Further, the computation of Smith form of $A,-p^{r}I]$ is also accomplished efficiently for known primes $p$ using $p$-adic arithmetic. This is a computational insight which reverses the usual computation of Bezout identity for the of $\gcd{a,p^{r}}$ by using Euclidean division. It shows that direct $p$-adic arithmetic requires only modulo $p$ computations while Euclidean division to find this GCD requires division with respect to multiple divisors and dividends.
@@
    Appendix A, treats the case of solving a linear system modulo $q^{r}$ where $q$ is the power $p^{d}$ of a prime denoting the $p$-adic byte representation of numbers modulo $q^{r}$. The resulting algorithm shows that the Smith form computation modulo $q^{r}$ can be carried out in efficient byte operations pre-defined in the computer.
@@
The above points briefly summarize the contributions of this paper. We focused mainly on the theory behind the computation and have omitted numerical illustrations, which are truly out of scope and would have required a much longer article. This paper is based on the arxiv version [15].
@@ 1.3 Efficient computation by reduction of Smith form modulo $p^{r}$
    Another important issue leading to the difficulty of solving modular liner system modulo $n$ arises when $n$ has divisors of the form $p^{r(p)}$ with $p$ prime and $r(p)>1$. For such cases solving the reduced linear systems $Ax=b\mbox{ \rm{mod} }p$ over prime fields $\mathbb{F}_{p}$ does not lead to a solution of the system modulo $n$. Hence in such cases, efficient (black box) linear algebra method of [3] cannot be used. Such cases of $r(p)>1$ the modular systems have to be necessarily addressed using the Smith form (or invariant factors) of the matrix $A,-p^{r(p)}I]$. Reduction of Smith form computation over moduli $p^{r}$ using parallel computation (and CRT) has surprisingly not appeared in the past literature.
@@ 2 Solution of the integral linear system in terms of the Smith form
where $f_{i}$ are unique positive integers such that $f_{1}|f_{2}$, $f_{2}|f_{3}$$f_{r-1}|f_{r}$.
@@ 2.1 Solution of the modular linear system
Then we can determine the solutions of the modular system by solving (4) using the Smith form of [A,nI][A,-nI].
@@
P[A,nI]Q=[S,0]P[A,-nI]Q=[S,0]
where $S$ is a $k\times k$ diagonal nonsingular matrix of invariant factors of [A,nI][A,-nI] and $[S,0]$ is a $k\times(k+l)$ matrix with $l$ zero columns added. Denote $\tilde{b}=Pb$. The system (4) has a solution iff $x_{0}=S^{-1}Pb$ is integral which is equivalent to $f_{i}|\beta_{i}$. For an arbitrary integral $l$ tuple $x_{1}$, denote $\tilde{x}=[x_{0}^{T},x_{1}^{T}]^{T}$. Then
@@
Since the modular system (1) has a solution iff the equivalent system (4) has an integral solution $x,y$, from the theorem (1) it follows that there exist unimodular matrices $P,Q$ which transform [A,nI][A,-nI] to $[S,0]$ since integral rank [A,nI][A,-nI] is equal to $k$. The matrix $S$ is the diagonal nonsingular matrix of invariant factors of [A,nI][A,-nI]. The integral system (4) is then transformed to
@@
for arbitrary $\mathbb{Z}^{l}$-tuple $x_{1}$. To express the conditions for uniqueness of solutions of the system (1) we need a definition of $\mbox{unimodular-}\mbox{rank}\,_{\mathbb{Z}_{n}}A$ of a matrix $A$ over $\mathbb{Z}_{n}$. Such a rank of $A$ is $r$ if $r$ is the largest size of minors of $A$ such that if $\{a_{1},a_{2},\ldots,a_{M}\}$ is the set of all such minors then the $GCD(a_{1},a_{2},\ldots,a_{M})$ is coprime to $n$.
@@
Assume first statement holds. Consider the residue equation $Ax=b\mbox{ \rm{mod} }p$ where $p$ is a prime divisor of $n$. If the system (1) has a unique solution then $x\mbox{ \rm{mod} }p$ is the unique solution of the residue equation at $p$ for all $p|n$. Hence $\mbox{rank}\,_{\mathbb{Z}_{p}}(A\mbox{ \rm{mod} }p)=r=l$ which must be less than or equal to $k$. This implies that for every prime divisor $p$ of $n$, there exists a non-zero $l\times l$ minor of $A$ modulo $p$. Since an $l\times l$ minor of $(A\mbox{ \rm{mod} }p)$ equals $l\times l$ minor of $A$ modulo $p$, it follows that the GCD of all $l\times l$ minors of $A$ is not divisible by $p$ for all prime divisors of $n$, i.e. the GCD of $l\times l$ minors of $A$ is coprime to $n$. Hence the condition that $\mbox{unimodular-}\mbox{rank}\,_{\mathbb{Z}_{n}}A=l$ follows. This proves that $1\Rightarrow 3$.
@@
$PAQ_{0}-nPQ_{2},-nPQ_{3}]\left[\begin{array}[]{l}x_{0}\\ x_{1}\end{array}\right]=Pb.$
@@ 2.2 Solution of the constrained system
Next consider the constraint $\phi(x)=\langle w,x\rangle$ required to be coprime to $n$ for some solution $x$ of (1) or equivalently that of (4). Consider the $(l+k)$ tuple $w^{T},0^{T}]^{T}$ obtained by adding $k$ tuple of zero rows to $w$. Denote
@@
where $Q$ is the unimodular matrix in the Smith form decomposition of [A,nI][A,-nI], $w_{0}$ is the first $l$ rows of $\tilde{w}$ and $w_{1}$ is the remaining $k$ rows of $\tilde{w}$. Then we have
@@
Let $\{p_{1},p_{2},\ldots,p_{r}\}$ denote the set of all prime divisors of $n$. Since all solutions of the modular system (1) are given by (6), the solutions which satisfy the constraint <w,x><w,x> coprime to $n$ need to be characterized within this set of all solutions. Existence of such solutions satisfying the constraint and choice of $x_{1}$ to construct a solutionare described by the next Theorem.
@@
[…]
@@
Theorem above shows how the Problem (1) can be solved by using the Smith form of [A,nI][A,-nI] and the knowledge of prime divisors of $n$.